Swifty Products, a rapidly growing distributor of home gardening equipment, is formulating its plans for the coming year. Carol Jones, the firm’s marketing director, has completed the following sales forecast.

Month Sales Month Sales
January $904,900 July $1,501,400
February $1,002,000 August $1,501,400
March $904,900 September $1,606,000
April $1,154,600 October $1,606,000
May $1,260,000 November $1,501,400
June $1,406,000 December $1,704,000


Phillip Smith, an accountant in the Planning and Budgeting Department, is responsible for preparing the cash flow projection. He has gathered the following information.

All sales are made on credit.
Swifty’s excellent record in accounts receivable collection is expected to continue, with 60% of billings collected in the month after sale and the remaining 40% collected two months after the sale.
Cost of goods sold, Swifty’s largest expense, is estimated to equal 40% of sales dollars. Seventy percent of inventory is purchased one month prior to sale and 30% during the month of sale. For example, in April, 30% of April cost of goods sold is purchased and 70% of May cost of goods sold is purchased.
All purchases are made on account. Historically, 75% of accounts payable have been paid during the month of purchase, and the remaining 25% in the month following purchase.
Hourly wages and fringe benefits, estimated at 30% of the current month’s sales, are paid in the month incurred.
General and administrative expenses are projected to be $1,575,000 for the year. A breakdown of the expenses follows. All expenditures are paid monthly throughout the year, with the exception of property taxes, which are paid in four equal installments at the end of each quarter.
   Salaries and fringe benefits $ 323,700
   Advertising 378,800
   Property taxes 142,400
   Insurance 195,600
   Utilities 184,800
   Depreciation 349,700
   Total $ 1,575,000
Operating income for the first quarter of the coming year is projected to be $329,100. Swifty is subject to a 40% tax rate. The company pays 100% of its estimated taxes in the month following the end of each quarter.

Swifty maintains a minimum cash balance of $50,000. If the cash balance is less than $50,000 at the end of the month, the company borrows against its 12% line of credit in order to maintain the balance. All borrowings are made at the beginning of the month, and all repayments are made at the end of the month (in increments of $1,000). Accrued interest is paid in full with each principal repayment. The projected cash balance on April 1 is $51,300.

Prepare the cash receipts budget for the second quarter. (Enter answers in necessary fields only. Leave other fields blank. Do not enter 0.)

Cash Receipts Budget

April

May

June

Total Cash Receipts

February sales

$Enter a dollar amount

$Enter a dollar amount

$Enter a dollar amount

$Enter a dollar amount

March sales

Enter a dollar amount

Enter a dollar amount

Enter a dollar amount

Enter a dollar amount

April sales

Enter a dollar amount

Enter a dollar amount

Enter a dollar amount

Enter a dollar amount

May sales

Enter a dollar amount

Enter a dollar amount

Enter a dollar amount

Enter a dollar amount

Totals $Enter a total amount

$Enter a total amount

$Enter a total amount

$Enter a total amount

Accounts Receivable balance at the end of second quarter of 2015 $Enter a dollar amount of the Accounts Receivable balance at the end of second quarter of 2015
Answer #1
April May June Total Cash Receipts
Februrary Sales $       4,00,800 $         4,00,800
March Sales $       5,42,940 $         3,61,960 $         9,04,900
April Sales $         6,92,760 $       4,61,840 $       11,54,600
May Sales $       7,56,000 $         7,56,000
Totals $       9,43,740 $       10,54,720 $    12,17,840 $       32,16,300
April May June
Februrary Sales =1002000*40%
March Sales =904900*60% =904900*40%
April Sales =1154600*60% =1154600*40%
May Sales =1260000*60%

Accounts Receivable Balance at end of second quarter = 40% of May Sales + Sales of June
= $1260000 x 40% + $1406000 = $1910000